OpenStudy (anonymous):

Find the arc length of the parametric equation : x= e^t , y= e^(-t) ,z= √2 t 0<=t<=1

OpenStudy (jamesj):

So you want\[\int\limits_0^1 \sqrt{ (x')^2 + (y')^2 + (z')^2 } dt\]

OpenStudy (anonymous):

yes, but i can't solve the equation...........please help !!!!

OpenStudy (jamesj):

Now the expression in the integral isn't as bad as it seems once you evaluate it. Hint: cosh^2 t - sinh^2 t = 1

OpenStudy (jamesj):

and remember also that you can write sinh t = 1/2(e^t - e^(-t))

OpenStudy (anonymous):

I'll try it out......thx for the hint !!

OpenStudy (jamesj):

sure. I can tell you the answer, but I think you'll enjoy seeing how this works out. It's quite elegant.

OpenStudy (jamesj):

Actually, you don't even need the hyperbolic functions if you can see the factorization, but I think it's helpful to know. The other one for the record is: cosh t = (1/2)(e^t + e^(-t))

OpenStudy (anonymous):

I obtain sth like this : and I wonder if it's wrong..... :(